### v1

# Learning What’s Missing: Attention Dispersion and EMA Stabilization in Length Generalization

Abstract
We study length generalization in transformers through the set complement task, where a model must predict a uniform distribution over tokens absent from an input sequence—an ability central to board-game style reasoning. Our main theoretical result establishes two statements. First, we prove tight bounds on embedding and value dimensions for single-layer attention-only transformers. Second, we show that if such a model achieves balanced logit displacement at lengths 1 and 2, then it must generalize to longer sequences, though with reduced precision. A mechanistic reading of the proof explains this limitation: as more tokens are attended to, softmax compresses logit displacements, eroding separation between valid and invalid outputs. Training dynamics also suggest a second obstacle: when many next tokens are possible, updates become noisy. We hypothesize that dropout can counteract the first effect and Exponential Moving Average (EMA) the second. We validate these hypotheses through random hyperparameter search on the set complement task, which confirms both mechanisms. We then test OthelloGPT, a GPT-1 style model trained on random Othello moves, and find that EMA again improves length generalization in this more complex setting.
1 Introduction
Since it was discovered that transformer [1]-based large language models (LLMs) can be aligned with human interests [2], LLM agents are being deployed in roles as diverse as application developers, counsellors, job interviewers, or research assistants. For safety and efficiency both it is paramount that we understand how these agents make their decisions.
Prior work revealed that to improve the reasoning capabilities of transformers, one can prompt [3] or train them to think in small steps. This makes reasoning transformers produce their answers akin to game-playing agents that map out various potential trajectories before making a move.
Therefore, we can gain insights on the reasoning processes of LLM agents by studying how models with similar architecture learn to play games. In the present work, we focus on the most fundamental skill an agent playing a game as simple as tic-tac-toe or as complex as go has to acquire: tell which board positions are not yet taken.
We abstract this task as the Set Complement Task, introduced in Subsection 3.1: given an input sequence of tokens without repetition, the model has to output a uniform distribution on the tokens absent from the input sequence. Note that we aim for more that top-1 accuracy, that is for the model to predict as most probable next token one that is missing from the input: such a basic component has to be learnt free from bias.
Our theoretical contribution is Theorem 4.2: a characterization of single-layer, attention-only, uniform attention models that can learn the task. First of all, we give tight bounds for the embedding and value dimensions required of the model. Second, we show that if the model can solve the task on input sequences of length 1 and 2, then it can solve the task for input sequences of any lengths, albeit at reduced precision.
This connects our work to the topic of length generalization: if a model robustly learned to perform an algorithmic task, then it should be able to produce a correct output on input sequences longer than those in its training set. It is an active field of study which tasks can transformers length generalize on: we discuss this in detail in Section 2.
It is of particular interest in the study of length generalization how to overcome obstacles to it. In Subsection 4.3, a mechanistic reading of the way our models make their inferences identifies the reduction in precision as a particular case of attention dispersion [4, 5]: softmax attention reduces displacements between attention weights as sequence length increases. We hypothesize that increased dropout may be able to increase the amplitude of value vectors and thus mitigate this effect.
We turn to the study of training dynamics in Subsection 4.5. We use conventional next token logit prediction training via negative log likelihood of one-hot sampled target distributions. Mechanistic analysis of training unveils a further obstacle for length generalization: in our task, the tokens that follow short sequences are sampled from many possibilities, which makes gradients noisy. Therefore, in Subsection 4.5, we make our second hypothesis: using the stabilizer Bias-corrected Exponential Moving Average (BEMA) may attenuate this effect.
We investigate the effect of our proposed strategies in a random hyperparameter search experiment, the results of which are reported in Section 5. We find that both our strategies reliably improve performance metrics.
To validate our methods in a more complex setting, in Section 6, we study length generalization in OthelloGPT [6, 7]. In this setup, we train GPT-1 style models to predict legal moves in random Othello games. We find that BEMA robustly improves performance metrics in this case too.
To summarize, our contributions are as follows:
  1. 1.
    We introduce the Set Complement Task, a simple algorithmic task that any board game playing agent should be able to solve and which can serve as a benchmark for length generalization and training with noisy gradients.
  2. 2.
    We characterize the minimal transformers that can solve the task. In particular, we prove a length generalization property: if a model can solve the task for input sequences of length 1 and 2 in a balanced way, then it must length generalize, albeit with reduced precision.
  3. 3.
    We mechanistically analyze inference and training dynamics, pointing out attention dispersion and noisy gradients as potential obstacles for length generalization. We propose increased dropout and BEMA as respective mitigations.
  4. 4.
    We conduct a random hyperparameter search experiment and show that our proposed strategies reliably improve performance metrics.
  5. 5.
    We study how our methods generalize to length generalization in OthelloGPT. We show that BEMA robustly improves performance in this setting too.
2 Related Work
Mechanistic Interpretability aims to find minimal subnetworks, so-called circuits, of an artificial neural network that satisfy a given task. In the case of transformers, several such circuits have been identified such as induction heads [8], indirect object identification circuits [9], greater-than circuits [10], and retrieval heads [11]. Of particular interest to the present work are the studies on OthelloGPT, which showed that in the residual stream of a GPT-1 style model trained to predict legal moves on random Othello games, via nonlinear [6] and linear [7] probing, one can find representations of board state. We intend to extend the compendium of known circuits by minimal transformers that can solve the Set Completion Task.
Length Generalization studies the conditions under which sequence-to-sequence models retain their performance on inputs longer than those seen during training. One train of results seeks to find criteria for algorithms transformers can length generalize on [12]. An important theoretical tool in this direction is the Restricted Access Sequence Processing Language (RASP) [13], a programming language that a transformer can implement. It was conjectured that length generalization is possible if there is a simple implementation in RASP-L [14]. Afterwards, a version of this conjecture was proven [15] using limit transformers, and a version of the C-RASP language [16]. In the usual algorithmic approach to the study of length generalization, if multiple solutions are possible, then the model has to predict the set of valid solutions as a singleton. We aim to bring in an alternative point of view closer to the spirit of language modeling: if multiple solutions are possible, the model should output a uniform distribution between them.
Attention Dispersion [4, 5] is a drawback of softmax attention that forbids generalization to arbitrary sequence lengths both in toy and language models. Most proposed solutions, such as adjusting attention logits [17] or next token logit temperatures [5] by a sequence length-dependent terms, make a change to the most commonly used architectures. In the present work, we are interested in whether we can use solutions that are still in line with mainstream models.
The study of the next token distributions output by LLMs brings a detailed view on how the models generate theirs answers, and how expressive they can get. An important part of this point of view is how calibrated are the models, that is how well do the predicted next token distributions approximate the target next token distribution. It is shown [18] that pretrained models as small as GPT-Neo-1.3B surpass humans in next token prediction on the OpenWebText dataset [19], both in top-1 accuracy, and perplexity. However, calibration to the pretraining corpus can be proven to bring in hallucinations [20], at the very least on facts not present in the training dataset, given the assumption that there are exponentially more ways to complete a sentence in an untruthful way. Neither base or aligned models are calibrated in numeric contexts such as generating tokens from a uniform distribution [21], rather they have strong systematic biases such as dependence on token order. Through soft [22] and hard [23] prompt tuning experiments, it was discovered [24] that transformers are more capable of outputting distributions of very low or very high entropy, those with outliers, or those that were output by other transformers. The experiments were conducted both on pretrained and randomly initialized models, thus indicating that the limits in expressivity may stem from the transformer architecture, or the softmax output. In our work, we also investigate if the model learns the uniform distribution among valid next tokens, thus indicating that predictions are free from bias.
3 Preliminaries
3.1 The Set Complement Task
In what follows, we shall introduce the set complement task that the models we interpret are trained on. To put it very succintly, the models are required to output a uniform distribution over tokens absent from an input without repetitions. Let us formalize this setting.
Let $v$ denote the number of distinct tokens. As they are only meant to signify the $v$ distinct elements of a finite set, we will denote tokens by integers. We let the vocabulary or ambient set be the finite set $\mathbb{V}=\{1,\dotsc,v\}$ of $v$ of distinct tokens. We call $v$ the vocabulary size or ambient set size. As it is not our focus here, we will forego using special tokens such as beginning of sequence, end of sequence, or padding. The valid input sequences are sequences $\mathbf{t}=(t_{1},\dotsc,t_{s})\in\mathbb{V}^{s}$ of length $1\leq s<v$ without repetitions: for distinct indices $1\leq i\neq j\leq s$, we have $t_{i}\neq t_{j}$. The underlying set of $\mathbf{t}$ is the set $\mathbf{t}]=\{t_{1},\dotsc,t_{s}\}$ of tokens in $\mathbf{t}$. We let $|\mathbf{t}|=s$ denote the length of $\mathbf{t}$.
We represent the set of categorical distributions on $v$ entries as the probability (v1)(v-1)-simplex
$\Delta^{v-1}=\left\{\mathbf{p}\in\mathbb{R}^{v}_{\geq 0}:\sum_{t=1}^{v}p_{t}=1\right\}.$
Let $\mathbb{X}$ denote the set of valid input sequences. Then the perfect solution to the task is the function $p^{*}:\mathbb{X}\to\Delta^{v-1}$ such that, for any input sequence $\mathbf{t}\in\mathbb{X}$ and token $t\in\mathbb{V}$, we have
$p^{*}(\mathbf{t})_{t}=\begin{cases}0&i\in[\mathbf{t}],\\ \frac{1}{v-s}&i\notin[\mathbf{t}].\end{cases}$ (1)
3.2 Minimal Transformers
We will seek to approximately solve the set complement task with parametric models of the form
$\mathbb{X}\xrightarrow{f_{\theta}}\mathbb{R}^{v}\xrightarrow{\mathrm{softmax}}\Delta^{v-1},$ (2)
where $f_{\theta}$ denotes a single-layer, attention-only, single-head, decoder-only transformer with parameter vector $\theta$. Let $\mathbf{t}=(t_{1},\dots,t_{s})\in\mathbb{X}$ be an input sequence. We call the output $f_{\theta}(\mathbf{t})\in\mathbb{R}^{v}$ the next token logit vector. The next token logit vector is formed as the linear combination
$f_{\theta}(\mathbf{t})=\mathbf{B}_{t_{s},:}+\sum_{i=1}^{s}\frac{a_{t_{s},t_{i}}}{\sum_{i^{\prime}=1}^{s}a_{t_{s},t_{i^{\prime}}}}\mathbf{D}_{t_{i},:},$ (3)
the terms in which are defined as follows:
The next token logit bias matrix $\mathbf{B}=\mathbf{E}\mathbf{U}\in\mathbb{R}^{v\times v}$ is the product of the token embedding parameter matrix $\mathbf{E}\in\mathbb{R}^{v\times d}$ and the unembedding parameter matrix $\mathbf{U}\in\mathbb{R}^{d\times v}$. For any $1\leq t\leq v$, the $t$-th row $\mathbf{E}_{t,:}\in\mathbb{R}^{d}$ is a $d$-dimensional trainable vector the token $t\in\mathbb{V}$ is mapped to. We call $d$ the embedding dimension, the vector space $\mathbb{R}^{d}$ the residual stream, and the vector space $\mathbb{R}^{v}$ the logit space. In our minimal transformers, we do not use positional encodings. During training, dropout [25] is optionally applied to the token embedding vectors. We call this dropout operation the embedding dropout.
The unnormalized per-token attention weight matrix $\mathbf{A}$ is formed as follows: First, RMSNorm [26] is applied to the token embedding vectors. We denote by $\mathbf{E}^{\prime}$ the token embedding parameter matrix with RMSNorm applied to its rows. Then, we form the query and key per-token matrices $\mathbf{Q}=\mathbf{E}^{\prime}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{E}^{\prime}\mathbf{W}_{K}$ via the query and key parameters matrices $\mathbf{W}_{Q},\mathbf{W}_{K}\in\mathbb{R}^{d\times d_{k}}$. We call $d_{k}$ the key dimension. In our setup, the key dimension is not necessarily equal to the embedding dimension divided by the number of attention heads. The per-token attention logit matrix is the product $\mathbf{A}^{\prime}=\mathbf{Q}\mathbf{K}^{T}$. This yields the unnormalized per-token attention weight matrix via the elementwise formula: $a_{i,j}=\exp(a^{\prime}_{i,j}/\sqrt{d_{k}})$. Note that as we define the output in Equation (3) for one input sequence only, we do not have to be explicit about causal attention. As we found it to be detrimental in initial experiments on the set complement task, we do not use attention dropout.
The next token logit displacement matrix $\mathbf{D}=\mathbf{E}^{\prime}\mathbf{W}_{V}\mathbf{W}_{O}\mathbf{U}$ is formed via the value and output parameter matrices $\mathbf{W}_{V}\in\mathbb{R}^{d\times d_{v}}$, $\mathbf{W}_{O}\in\mathbb{R}^{d_{v}\times d}$. We call $d_{v}$ the value dimension. In our setup, the value dimension is not necessarily equal to the key dimension, nor is it necessarily equal to the embedding dimension divided by the number of attention heads. During training, dropout is optionally applied to the rows of the matrix product $\mathbf{E}^{\prime}\mathbf{W}_{V}\mathbf{W}_{O}$ before multiplication from the right by the unembedding parameter matrix $\mathbf{U}$. We call this dropout operation the residual dropout.
4 Theoretical Analysis
We say that the model $f_{\theta}\colon\mathbb{X}\to\mathbb{R}^{v}$ has precision $C>0$ if for all input sequences $\mathbf{t}\in\mathbb{X}$, and tokens $u\in\mathbb{V}\setminus[\mathbf{t}]$, $v\in\mathbb{V}$, we have
$f_{\theta}(\mathbf{t})_{u}-f_{\theta}(\mathbf{t})_{v}\begin{cases}>C&v\in[\mathbf{t}],\\ =0&v\in\mathbb{X}\setminus[\mathbf{t}].\end{cases}$ (4)
We say that the model has precision $C>0$ at (resp. up to) length $s$, if the above property (4) is satisfied for input sequences $\mathbf{t}\in\mathbb{X}$ of length $|\mathbf{t}|=s$ (resp. $\leq s$).
4.1 A Hardcoded, Minimal Solution
Let us first provide a hardcoded model that is precise up to level $D$. In Theorem 4.2, we will show that its embedding and key dimensions v1v-1 are actually the smallest possible dimensions with which it is possible to solve the task.
Example 4.1.
For any vocabulary size $v$, we now give a formula for a model that is arbitrarily close to being perfect if we choose $C>0$ large enough. It uses embedding and value dimensions $d=d_{v}=v-1$, and key dimension $d_{k}=1$. We can use the following parameter matrices:
$\mathbf{E}=\begin{pmatrix}1&0&\dotsb&0\\ 0&1&\dotsb&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dotsb&1\\ -1&-1&\dotsb&-1\end{pmatrix},\,\mathbf{U}=\begin{pmatrix}-1&0&\dotsb&0&0\\ 0&-1&\dotsb&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dotsb&-1&0\end{pmatrix}$,
$\mathbf{W}_{Q}=\mathbf{W}_{K}=\boldsymbol{0},\,\mathbf{W}_{V}=vC\mathbf{I},\,\mathbf{W}_{O}=I.$
As the embedding vectors have constant length, we can set the RMSNorm scaling parameters to get $\mathbf{E}^{\prime}=\mathbf{E}$. Note that with these parameters, we get
$\mathbf{B}=\begin{pmatrix}-1&0&\dotsb&0&0\\ 0&-1&\dotsb&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dotsb&-1&0\\ 1&1&\dotsb&1&0\end{pmatrix},\,\mathbf{A}=\boldsymbol{1},\,$
$\mathbf{D}=\begin{pmatrix}-vC&0&\dotsb&0&0\\ 0&-vC&\dotsb&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dotsb&-vC&0\\ vC&vC&\dotsb&vC&0\end{pmatrix}$.
Thus, one can check by hand that $f_{\theta}$ has precision $C$.
Note that the hardcoded model $f_{\theta}$ has constant attention. Since when training transformers with weight decay induces low-rank attention logit matrices [27], we will continue our theoretical investigation with the assumption of constant attention. Therefore, the formula (3) for the next token logit vector $f_{\theta}(\mathbf{t})$ at the input sequence $\mathbf{t}=(t_{1},\dotsc,t_{s})\in\mathbb{X}$ simplifies to the following:
$f_{\theta}(\mathbf{t})=\mathbf{B}_{t_{s},:}+\frac{1}{s}\sum_{i=1}^{s}\mathbf{D}_{t_{i},:}.$ (5)
4.2 Length Generalization at the Price of Less Precision
In this Subsection, we prove tight bounds on the embedding and value dimensions of a constant attention model $f_{\theta}$. Moreover, we show that if $f_{\theta}$ approximates the ideal solution on lengths 1 and 2, and moreover it satisfies a balance criterion on token displacements, then it length generalizates, albeit with decreasing precision as length increases.
Theorem 4.2.
Assume that the model $f_{\theta}$ has constant attention. Then the following statements hold:
(a) Suppose that the model $f_{\theta}$ has precision $C>0$ at length 1. Then the matrix $\mathbf{B}+\mathbf{D}$ has rank at least v1v-1. In particular, we have $d\geq v-1$.
(b) Suppose moreover that the model $f_{\theta}$ also has precision $C>0$ at length 2. Then the matrix $\mathbf{D}$ also has rank at least v1v-1. In particular, we have $d_{v}\geq v-1$.
(c) Suppose moreover that the following condition is satisfied: for all pairs of distinct tokens $t,u\in\mathbb{V}$, we have
$f_{\theta}((t))_{u}-f_{\theta}((t))_{t}<2C.$ (6)
Then for each $3\leq s<v$, the model $f_{\theta}$ has precision $\frac{2}{s}C$ at length $s$.
The following Lemma is a key component of the tight dimension bounds:
Lemma 4.3.
Let $\mathbf{u},\mathbf{v},\mathbf{w}\in\mathbb{R}^{n}$ be $n$-dimensional vectors. Suppose that we have $w_{i}<0$ for all indices $1\leq i\leq n$. Then the matrix $\mathbf{A}:=\boldsymbol{1}\mathbf{u}^{T}+\mathbf{v}\boldsymbol{1}^{T}+\mathop{\mathrm{diag}}(\mathbf{w})$ has rank at least n1n-1.
Proof.
It is enough to show that the matrix $\mathbf{A}$ is injective on the 1-codimensional subspace $Z:=\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}=0\}$. Take $\mathbf{x}\in Z$ and suppose that we have $\mathbf{A}\mathbf{x}=\boldsymbol{0}$. Let $\alpha=\mathbf{u}^{T}\mathbf{x}$. Then for each $1\leq i\leq n$, we have
$0=(\mathbf{A}\mathbf{x})_{i}=\alpha+v_{i}\sum_{i=1}^{n}x_{i}+w_{i}x_{i}=\alpha+w_{i}x_{i}.$
If $\alpha=0$, then as $w_{i}<0$ for all $1\leq i\leq n$, we get $\mathbf{x}=\boldsymbol{0}$. Otherwise, we get
$0=\sum_{i=1}^{n}x_{i}=-\alpha\sum_{i=1}^{n}\frac{1}{w_{i}}<0,$
a contradiction.
Proof of Theorem 4.2.
(a) In terms of the matrices $\mathbf{B}$ and $\mathbf{D}$, the fact that the model $f_{\theta}$ has precision $C$ at length 1 reads as, for distinct tokens $t,u,v\in\mathbb{V}$:
$b_{t,u}+d_{t,u}$ $>b_{t,t}+d_{t,t}+C$ (7)
$b_{t,u}+d_{t,u}$ $=b_{t,v}+d_{t,v}$ (8)
These conditions imply the conditions of Lemma 4.3 for the matrix $\mathbf{B}+\mathbf{D}$, thus showing that we have $\mathop{\mathrm{rank}}(\mathbf{B}+\mathbf{D})\geq v-1$.
(b) In terms of the matrices $\mathbf{B}$ and $\mathbf{D}$, the fact that the model $f_{\theta}$ has precision $C$ at length 2 reads as, for distinct tokens $t,u,v,w\in\mathbb{V}$:
$2b_{t,v}+d_{t,v}+d_{u,v}$ $>2b_{t,t}+d_{t,t}+d_{u,t}+2C$ (9)
$2b_{t,v}+d_{t,v}+d_{u,v}$ $>2b_{t,u}+d_{t,u}+d_{u,u}+2C$ (10)
$2b_{t,v}+d_{t,v}+d_{u,v}$ $=2b_{t,w}+d_{t,w}+d_{u,w}.$ (11)
Equations (8) and (11) show that for all distinct tokens $t,u,v\in\mathbb{V}$: the difference $d_{t,v}-d_{u,v}$ is constant in $v$. Let us denote this by $\alpha_{t,u}$, and let $\alpha_{t,t}:=0$.
Let us fix $r\in\mathbb{V}$ and let $\mathbf{a},\mathbf{c}\in\mathbb{R}^{v}$ be defined by $a_{t}=\alpha_{t,r},c_{t}=d_{r,t}$ for $t\in\mathbb{V}$. Then for all distinct $t,u\in\mathbb{V}$: we have $d_{t,u}=a_{t}+c_{u}$. Moreover, Constraints (8) and (10) show that we have $d_{t,t}-a_{t}-c_{t}<0$. Therefore, Lemma 4.3 shows that we have $\mathop{\mathrm{rank}}(\mathbf{D})\geq v-1$.
(c) Let us prove that $f_{\theta}$ has precision $\frac{2}{s}C>0$ at length $s$ by induction on $1\leq s<v$. By assumption, the induction hypothesis holds for $s=1,2$. Let us assume that it holds for $s$, that is, the following constraints are satisfied, for distinct tokens $t_{1},\dotsc,t_{s},u,v\in\mathbb{V}$, and indices $1\leq i<s$:
$sb_{t_{s},u}+d_{t_{1},u}+\dotsb+d_{t_{s},u}$ $>sb_{t_{s},t_{s}}+d_{t_{1},t_{s}}+\dotsb+d_{t_{s},t_{s}}+2C$ (12)
$sb_{t_{s},u}+d_{t_{1},u}+\dotsb+d_{t_{s},u}$ $>sb_{t_{s},t_{i}}+d_{t_{1},t_{i}}+\dotsb+d_{t_{s},t_{i}}+2C$ (13)
$sb_{t_{s},u}+d_{t_{1},u}+\dotsb+d_{t_{s},u}$ $=sb_{t_{s},w}+d_{t_{1},w}+\dotsb+d_{t_{s},w}.$ (14)
Let us undertake proving the induction step. By Inequalities (9) and (12), we get
$(2b_{t_{s+1},u}+d_{t_{1},u}+d_{t_{s+1},u})+(sb_{t_{s+1},u}+d_{t_{2},u}+\dotsb+d_{t_{s+1},u})$
$>$ $(2b_{t_{s+1},t_{s+1}}+d_{t_{1},t_{s+1}}+d_{t_{s+1},t_{s+1}})$
$+(sb_{t_{s+1},t_{s+1}}+d_{t_{2},t_{s+1}}+\dotsb+d_{t_{s+1},t_{s+1}})+2C+2C,$
which by Inequality (6) yields
$(s+1)b_{t_{s+1},u}+d_{t_{1},u}+\dotsb+d_{t_{s+1},u}$
$>$ $(s+1)b_{t_{s+1},t_{s+1}}+d_{t_{1},t_{s+1}}+\dotsb+d_{t_{s+1},t_{s+1}}+2C.$
Then note that by Equation (8), Equation (11) is equivalent to the following equation:
$b_{t_{s+1},u}+d_{1,u}=d_{t_{s+1},w}+d_{1,w}.$ (15)
With this and Inequality (13), we get
$(b_{t_{s+1},u}+d_{1,u}+d_{t_{s+1},u})+(sb_{t_{s+1},u}+d_{t_{2},u}+\dotsb+d_{t_{s+1},u})$
$>$ $(b_{t_{s+1},t_{j}}+d_{1,u}+d_{t_{s+1},t_{j}})+(sb_{t_{s+1},t_{j}}+d_{t_{2},u}+\dotsb+d_{t_{s+1},t_{j}})+2C$
Finally, Equalities (15) and (14) yield
$(b_{t_{s+1},u}+d_{1,u}+d_{t_{s+1},u})+(sb_{t_{s+1},u}+d_{t_{2},u}+\dotsb+d_{t_{s+1},u})$
$=$ $(b_{t_{s+1},v}+d_{1,u}+d_{t_{s+1},v})+(sb_{t_{s+1},v}+d_{t_{2},u}+\dotsb+d_{t_{s+1},v})$
Figure 1: Summary of best metrics per model. We do not display ITR values as they are below $5\cdot 10^{-4}$ in all cases.
4.3 Resolving Attention Dispersion by Dropout
Table 1: Hyperparameter distributions for the random search
Hyperparameter Distribution Range
Dataloader
input sequence size $s$ $1+\lfloor 2^{\mathcal{U}[0,4]}\rfloor$ $[2,17]$
ambient set size $v$ $s+\lfloor 2^{\mathcal{U}[0,4]}\rfloor$ $[3,33]$
Model
embedding dim $d$ $\lfloor(v-1)\mathcal{U}[1,4]\rfloor$ $[2,128]$
key dim $d_{k}$ $\lfloor\mathcal{U}[1,d]\rfloor$ $[1,128]$
value dim $d_{v}$ $\lfloor\mathcal{U}[v-1,d]\rfloor$ $[2,128]$
RMSNorm $\epsilon$ $^{\mathcal{U}[-10,-4]} [1010$$^{-10},10^{-4}$]
AdamW
1. moment decay $\beta_{1}$ $1-10^{\mathcal{U}[-2,0]}$ $0,1-10^{-2}]$
2. moment decay $\beta_{2}$ $10^{\mathcal{U}[-1,-8]} [11$$10^{-1},1-10^{-8}$]
weight decay $\lambda$ $10^{\mathcal{U}[-6,0]}$ $10^{-6},1]$
AdamW $\epsilon$ $10^{\mathcal{U}[-12,-8]}$ $10^{-12},10^{-8}]$
max gradient norm $10^{\mathcal{U}[-2,2]}$ $10^{-2},10^{2}]$
Learning Rate Schedule
peak learning rate $\eta$ $10^{\mathcal{U}[-5,-1]}$ $10^{-5},10^{-1}]$
warmup steps $\lfloor 10^{\mathcal{U}[-2,6]}\rfloor$ $0,10^{6}]$
multiplier at end $10^{\mathcal{U}[-4,0]}$ $10^{-4},1]$
Dropout
embedding dropout $\mathop{\mathrm{relu}}(\mathcal{U}[-\frac{1}{2},\frac{1}{2}])$ $0,\frac{1}{2}]$
residual dropout $\mathop{\mathrm{relu}}(\mathcal{U}[-\frac{1}{2},\frac{1}{2}])$ $0,\frac{1}{2}]$
BEMA
BEMA power $\eta$ $\mathcal{U}[0,1]$ $[0,1]$
EMA lag $\rho$ $0^{\mathcal{U}[0,10]} [1,1$$10^{10}$]
EMA power $\kappa$ $\mathcal{U}[0,1]$ $[0,1]$
Inspection of formula (5) shows how precision decreases with length: even if parameters are learnt that output precise results on small sequences as
$f_{\theta}((t_{1}))=\mathbf{B}_{t_{1},:}+\mathbf{D}_{t_{1},:}\text{ and }f_{\theta}((t_{1},t_{2}))=\mathbf{B}_{t_{2},:}+\frac{\mathbf{D}_{t_{1},:}+\mathbf{D}_{t_{2},:}}{2},$
in longer sequences, softmax attention dilutes the next token logit displacements $\mathbf{D}$.
To resolve this dispersion problem, we want the model to learn larger next token logit displacements, even in the presence of weight decay. We hypothesize that increased dropout can have this effect: in training, random subnetworks learn smaller displacements that are enough to be precise on shorter sequences. Then, during inference, the smaller displacements are accumulated into larger ones, thus counteracting the dispersion effect.
4.4 Training: NLL of One-Hot Sampled Target Distribution
We seek to get models with next token probability distribution $p_{\theta}(\mathbf{t})=\mathop{\mathrm{softmax}}(f_{\theta}(\mathbf{t}))\in\Delta^{v-1}$ approximating the uniform distribution $p^{*}(\mathbf{t})$ on tokens absent from the input sequence $\mathbf{t}=(t_{1},\dotsc,t_{s})\in\mathbb{X}$, see Equation (1). However, in our study of training dynamics, we intend to follow the general practice in training generative language models: we sample an extra token $u\in\mathbb{V}\backslash[\mathbf{t}]$ and the model receives as loss the negative log likelihood
$\mathop{\mathrm{NLL}}(\mathbf{t},u;\theta)=-\log p_{\theta}(\mathbf{t})_{u}$
between the predicted next token logit distributions after the input sequence $\mathbf{t}$ and the one-hot categorical distribution at token $u$.
As we have $u\notin[\mathbf{t}]$, the concatenation $(t_{1},\dotsc,t_{s},u)\in\mathbb{V}^{s+1}$ has no repetitions and is thus a valid input sequence if and only if we have s<v1s<v-1. We let $\bar{\mathbb{X}}$ denote the collection of sequences of tokens from $\mathbb{V}$ without repetition. That is, for a sequence $\mathbf{t}\in\bar{\mathbb{X}}$, we have $\mathbf{t}\in\mathbb{X}$ if and only if $|\mathbf{t}|<v$. For a prefix length $1\leq s^{\prime}\leq|\mathbf{t}|$, the prefix sequence of length $s^{\prime}$ is $\mathbf{t}_{:s^{\prime}}=(t_{1},\dotsc,t_{s^{\prime}})\in\bar{\mathbb{X}}$.
In our case of main interest, that of length generalization, the lengths $s$ of input sequences are much smaller than the ambient set size $v$. This means that the number vsv-s of possible target next tokens $u\in\mathbb{V}\backslash[\mathbf{t}]$ is large. Therefore, the model will receive training signals with high noise, thus slowing training.
4.5 Mitigating Slowdown from Noisy Sampled Targets with BEMA
Figure 2: Mean dropout rates of top portions of models per generalization gap and extra dimensions.
We hypothesize that Exponential Moving Average (EMA), a general remedy for gradient noise-induced slowdown, may provide a mitigation in this setting. We use Bias-Corrected Exponential Moving Average (BEMA) [28], that we now introduce for completeness:
We use three hyperparameters: the EMA lag $\rho$, the EMA power $\kappa$, and the BEMA power $\eta$. At training step $n\geq 0$, let $\theta_{n}$ denote the parameter values. In particular, we denote the intial parameter values as $\theta_{0}$. In the context of EMA, we also call them training parameter values. Then the EMA parameter values $\theta^{\mathrm{EMA}}_{n}$ are inductively defined as follows:
$\theta^{\mathrm{EMA}}_{0}$ $=\theta_{0}\text{ and }$
$\theta^{\mathrm{EMA}}_{n+1}$ $=(1-\beta_{n})\theta^{\mathrm{EMA}}_{n}+\beta_{n}\theta_{n+1}\text{ where }\beta_{n}=(\rho+n)^{-\kappa}.$
We call $\beta_{n}$ an EMA weight. Finally, at inference, we use the BEMA parameter values $\theta^{\mathrm{BEMA}}_{n}$, that are defined as follows:
$\theta^{\mathrm{BEMA}}_{n}=\alpha_{n}(\theta_{n}-\theta_{0})+\theta^{\mathrm{EMA}}_{n}\text{ where }\alpha_{n}=(\rho+n)^{-\eta}.$
We call $\alpha_{n}$ a BEMA weight.
Note that we only need to store, in addition to the most recent training parameter values $\theta_{n}$, the initial parameter values $\theta_{0}$, and the most recent EMA parameter values $\theta^{\mathrm{EMA}}_{n}$.
5 Random Search Experiments
5.1 Hyperparameter Distributions
We run a hyperparameter random search to see how different hyperparameter configurations influence performance, in particular length generalization. See Table 1 for the distributions we sample the hyperparameters from. We randomly sample 260 collections of dataloader hyperparameters $v,s$ and the archictectural hyperparameters $d,d_{k},d_{v}$. Then for each of these hyperparameter collection, we sample 1000 collections of hyperparameters such as learning rate or number of warmup steps that we can vectorize over. In particular, for each of the 1000 ensemble members, we sample a single collection $\rho,\kappa,\eta$ of BEMA hyperparameters.
5.2 Metrics
We introduce three metrics to measure how well a model $f_{\theta}$ solves the set complement task. Let $\mathbf{t}\in\mathbb{X}$ be an input sequence. Recall that $f_{\theta}(\mathbf{t})\in\mathbb{R}^{v}$ is the vector of predicted unnormalized next token logits to follow $\mathbf{t}$, and we let $p_{\theta}(\mathbf{t})=\mathop{\mathrm{softmax}}(f_{\theta}(\mathbf{t}))$ denote the corresponding next token probabilities.
The most important metric which we utilize to measure how closely the predicted distribution $p_{\theta}(\mathbf{t})$ approximates the uniform distribution on legal tokens $p^{*}(\mathbf{t})$ (see Equation (1)) is total variation distance (TVD):
$\mathop{\mathrm{TVD}}(\mathbf{t};\theta)=\frac{1}{2}\sum_{i=1}^{v}|p_{\theta}(\mathbf{t})_{i}-p^{*}(\mathbf{t})_{i}|.$ (16)
A more permissive metric that does not require uniformity on legal token probabilities is illegal token probability (ITP), that measures the total probability mass put on illegal tokens:
$\mathop{\mathrm{ITP}}(\mathbf{t};\theta)=\sum_{t\in[\mathbf{t}]}p_{\theta}(\mathbf{t})_{t}.$ (17)
A yet more permissive metric that only requires that the token with the highest probability—or logit—is legal is the illegal token rate (ITR):
$\mathop{\mathrm{ITR}}(\mathbf{t};\theta)=\boldsymbol{1}_{[\mathbf{t}]}(\mathop{\mathrm{argmax}}f_{\theta}(\mathbf{t}))=\begin{cases}1&\mathop{\mathrm{argmax}}(f_{\theta}(\mathbf{t}))\in[\mathbf{t}],\\ 0&\text{else}.\end{cases}$ (18)
5.3 Dataloaders, Loss and Metric Aggregation
Figure 3: Means of top quantiles of BEMA model metrics and no BEMA metrics in OthelloGPT length generalization.
Both our training and validation dataloaders output minibatches of sequences $\mathbf{T}=(\mathbf{t}_{1},\dotsc,\mathbf{t}_{N})\in\bar{\mathbb{X}}^{N}$. The difference is in the length of the sequences: in each of our training runs, we sample a training input sequence size $2\leq s<v$, and let the training dataloader output minibatches of sequences of length $s+1$. On the other hand, in each of our runs, we let the validation dataloader output minibatches of sequences of length v1v-1. In both cases, our dataloaders sample sequences uniformly.
Let $\mathbf{T}$ denote a training or validation minibatch. We follow the standard convention to aggregate the loss and the metrics not only by averaging across the minibatch entries, but also across the prefixes of the sequences. That is, the training loss on one minibatch is
$\mathop{\mathrm{NLL}}(\mathbf{T};\theta)=\frac{1}{Ns}\sum_{i=1}^{N}\sum_{s^{\prime}=1}^{s}\mathop{\mathrm{NLL}}(\mathbf{T}_{i,:s^{\prime}},t_{i,s^{\prime}};\theta)$
and if $\mu$ is one of the metrics defined in Subsection 5.2, then the corresponding validation metric reported in our experiments is
$\mu(\mathbf{T};\theta)=\frac{1}{Ns}\sum_{i=1}^{N}\sum_{s^{\prime}=1}^{s}\mu(\mathbf{T}_{i,:s^{\prime}};\theta).$
In training, we use minibatch size $N=128$, and we report validation results based on samples of $N=1024$ sequences.
5.4 Training
Following standard conventions, we initialize parameter matrices with normal distribution of std $\sigma=0.02$ and truncated at $2\sigma$. We train each ensemble of 1000 models for a maximum number of 10 000 000 AdamW [29] steps. We follow the standard practice of disabling weight decay on embedding, and norm vectors; the latter decision is ablated in [30]. We calculate validation metrics every 10 000 training steps. If there is no improvement in validation TVD of any ensemble member for 1 000 000 steps, then we stop early. For learning rate schedule, we use linear warmup, and linear decay [31]. For each of our models, we report two sets of metrics: one for the training parameters $\theta$, and one for the BEMA parameters $\theta^{\mathrm{BEMA}}$, see Subsection 4.5.
Training 260 000 models takes 1024 NVIDIA H100 NVL (96 GB) hours. We release our code upon publication, to support reproducibility, and further research.
5.5 Experiment Results
In reporting our results, we call the difference v1sv-1-s the generalization gap. Note that this quantity signifies the amount of length generalization the model has to perform: as specified in Subsection 5.3, the training input sequences have length $s$, while the validation input sequences have length v1v-1. To prove robustness of our results to hyperparameter choices, besides reporting the top metrics per model ensemble, we also report the mean metrics of the best 1% and 10%.
See Figure 1 for a summary of best metrics per model. One can see that BEMA indeed improves length generalization. Note that it does so robustly to the BEMA hyperparameters $\eta,\rho,\kappa$ (See Subsection 4.5 for definitions), as in each model we only sample a single set of BEMA hyperparameters. We do not display ITR values as they are below $5\cdot 10^{-4}$ in all cases: it is a much easier task to make a legal next token have the highest logit than to learn to output a uniform distribution on all legal next tokens.
Let us now turn to a demonstration of how dropout can boost performance across a generalization gap—if there are enough neurons so that the task can be learned with part of them shut down. Recall that (i) as per Theorem 4.2ab, the embedding dimension $d$ and value dimension $d_{v}$ both have to be at least v1v-1, and (ii) as written in Table 1, the embedding dimension $d$ is sampled from $\lfloor(v-1)\mathcal{U}[1,4]\rfloor$, and the value dimension $d_{v}$ is sampled from $\lfloor\mathcal{U}[v-1,d]\rfloor$. To display the number of extra parameters, we use the embedding dimension multiplier $\frac{d}{v-1}$, and the value dimension coefficient, that is 0 if d=v1d=v-1, and $\frac{d_{v}-v+1}{d-v+1}$ otherwise.
In Figure 2, we plot the average embedding and residual dropout rates, for the top 0.1%, 1%, and 10% of the models with and without BEMA. We see that, under substantial generalization gap, and if the embedding dimension multiplier is big enough, then better models often have increased embedding dimension. We see a less pronounced effect in the case of residual dropout.
6 Length Generalization in OthelloGPT
Figure 4: Metrics per BEMA power–EMA power pairs in OthelloGPT length generalization.
6.1 Experiment Setup
Table 2: Hyperparameter distributions for OthelloGPT
Hyperparameter Distribution
input sequence length $s$ $\lfloor\mathcal{U}[15,60)\rfloor$
attention dropout $\mathcal{U}[0,0.2]$
embedding dropout $\mathcal{U}[0,0.5]$
residual dropout $\mathcal{U}[0,0.3]$
To test how our findings generalize to a more complex setting, we perform length generalization experiments on training a GPT-1 [32] style model to output legal Othello moves via next token prediction on random Othello games. Previously, such models were studied to see if they learn a world model: the internal representations produced by such models were probed for nonlinear [6] and linear [7] representations of board state.
Besides the fixed 4 starting positions, an Othello game has $v=60$ positions. We test how well can a model trained on the first $15\leq s<60$ moves in a random game can output a uniform distribution on the next legal tokens. To measure the performance of the models, we use the same metrics as introduced in Subsection 5.3.
The original models use learned absolute positional embedding. Since that is incompatible with length generalization, we replace it by rotary positional encoding [33]. We also keep track of 10 EMA parameters $\theta^{\mathrm{EMA}}$, as defined in Subsection 4.5, with EMA power $\kappa=0.1,0.2,\dotsc,1$ and EMA lag $\rho=10$. At validation, we produce from each set of EMA parameters 10 BEMA parameters $\theta^{\mathrm{BEMA}}$ with BEMA power $\eta=0.1,0.2,\dotsc,1$. Thus, in total, we test 100 BEMA hyperparameter sets.
We sample 100 sets of input sequence length $s$ and dropout rates following Table 2. We train each model on a training set of 10 000 000 random Othello games for 1 epoch, validating every 1000 training steps. We use training minibatch size 256.
6.2 Experiment Results
Figure 5: Generalization gap to TVD without BEMA in OthelloGPT, colored by embedding dropout rate.
First of all, let’s see a summary of metrics per generalization gap in Figure 3. To indicate the robustness of the improvement BEMA brings, out of the 100 BEMA metrics, we report the best of 100, the mean of the top 10, and the mean of all the values. We can see that already the mean of the top 10 brings substantial improvement.
If we turn to metrics by BEMA hyperparameters in Figure 4, we see that the fact that in Figure 3 the mean metric over all BEMA hyperparameters was worse than the metric without BEMA was caused by the significantly worse results for EMA power values $\kappa=0.1,0.2$. In fact, BEMA performance is robust with regards hyperparameters, yielding good results for the range $0.4\leq\kappa\leq 0.9$, $0.5\leq\eta\leq 0.8$.
It turns out that in length generalization with OthelloGPT, dropout does not give a significant improvement. See for example Figure 5 for data on TVD without BEMA, and embedding dimension. It is up to further research to determine if the embedding or value dimensions were simply not big enough, or there is another effect at play. Note also the interesting detail that for generalization gap larger than 20, the gap to TVD values seem to follow a power law. This also merits futher investigation.
7 Conclusion
We introduce the Set Complement Task, that abstracts the fundamental skill of board game playing agents of detecting which positions are not yet taken. We prove that the minimal transformer models that can solve the task have to length generalize, albeit at the cost of reduced precision. Via mechanistic analysis, we uncover methods that can help mitigate said reduction. We show via random hyperparameter search that our methods are indeed effective. Finally, we show that BEMA helps length generalization in the more complex setting of OthelloGPT too.
Acknowledgments
Part of this research was conducted under the auspices of the Budapest Semesters in Mathematics program’s ”Research Opportunities” initiative. P. Zs. was supported by the Ministry of Innovation and Technology NRDI Office within the framework of the Artificial Intelligence National Laboratory (RRF- 2.3.1-21-2022-00004).